# Geometric insights into robust portfolio construction

### v4

@@ 1 Introduction
Using a geometric construction that relates the angle between the vector of expected returns $\boldsymbol{\alpha}$ and portfolio weights $\boldsymbol{\omega}$ determined under an unconstrained mean-variance optimisation, Golts and Jones [19] were able to quantify the distortions induced by the optimisation process without making direct distributional assumptions about the underlying data generating process. They were able to link this angle directly to the condition number of the covariance matrix by showing that this angle has an upper bound dependent on the condition number. The interpretation is novel because of its geometric construction and the weak distributional assumptions this allows. When a covariance matrix is ill-conditioned, or close to degenerate, the $\boldsymbol{\alpha}$-weight angle should be very large; whereas well-conditioned matrices ensure the angle remains within tight bounds. This is much more than merely a representation change, because it provides a geometric explanation why an equally weighted portfolio is not always preferable to a mean-variance solution even in the presence of estimation uncertainty.
@@
Reducing the angle between the weights and expected return vector is not equivalent to the implementation of a return maximisation, or return maximisation constraint. The case of pure return maximisation would result in a portfolio that weighs toward the asset with the greatest expected return, and not the direction of the expected return vector in general. In practice the $\boldsymbol{\alpha}$-weight angle should never be zero as there is always a residual correlation between the assets. In mean variance optimisations there is always a conundrum due to correlation specification error – and then determining the degree of shrinkage that should be adopted to best reflect this. The test of how much one should intervene can only be determined by out-of-period portfolio performance simulation, or with direct distributional assumptions (as demonstrated with simply simulations in Section 5).
@@ 1.1 Estimation uncertainty
Additionally, even if means and variances could accurately represent asset class distributions, the theory assumes both future expected returns, and the covariance matrix of asset returns are known, or even that they can be known. The presence of estimation error in input parameters unfortunately makes the use of Modern Portfolio Theory in its classical Markowitz formulation impractical – this is well understood and common cause. The short-comings of mean-variance optimisation has been extensively discussed in the literature, with highlights including the work by Best and Grauer [6], Frankfurter et al. [16], Jobson and Korkie [29, 30], Zimmermann and Nierdermayer [44], Broadie [8], Michaud [38], and Chopra and Ziemba [12].
@@
In order to ameliorate the effect of estimation errors in the estimates of expected returns, attempts have been made to create better and more stable mean-variance optimal portfolios by utilising expected return estimators that have a better behaviour when used in the context of the mean-variance framework. One of the more common techniques is the utilisation of James-Stein estimators (see Jobson and Korkie [30]) to shrink the expected returns towards the average expected return based on the volatility of the asset and the distance of its expected return from the average. Jorion [31] developed a similar technique that shrinks the expected return estimate towards the minimum variance portfolio. The area of robust statistics (see [10]) has recently been employed to create stable expected return estimates as well. Chopra and Ziemba [12] show that errors in the sample mean estimate have a larger impact on the out-of-sample performance than errors in the sample covariance estimate.
@@
Here we demonstrate this relationship geometrically and mathematically using an extension of the Kantorovich inequality [24, 22] provided by Bauer and Householder [4]; this elegant machinery will be shown as useful in determining why this relationship exists (outlines to both theorems are provided in A and B). Inspiration for this idea can be seen in [2] where much of the geometric interpretation around matrices is described.
@@ 1.2 The efficient frontier and $\alpha$-weight angle
Consider a portfolio with $n$ assets, expected return vector $\boldsymbol{\alpha}$, covariance matrix $\Sigma$, and a fixed portfolio mean $\alpha_{p}$. To find this portfolio on the efficient frontier with asset weights $\boldsymbol{\theta}$, we perform a portfolio variance minimization ([36, 37]):
@@
Table 1: Some classical unconstrained portfolio optimisation problems from [19]. Here portfolios employ no explicit gearing constraint, and positions may be unconstrained long/short. Here I is a return maximisation problem with risk bound, II is a risk minimisation problem with a return bound, III is a conventional mean-variance optimisation, and IV is the Sharpe ratio optimisation. In III $\gamma$ is the inverse risk aversion parameter. In IV the risk free rate $r_{f}$ is assumed to be zero. Here $A={\mathbf{1}^{\prime}\Sigma^{-1}\mathbf{1}}$,$B={\boldsymbol{\alpha}^{\prime}\Sigma^{-1}\mathbf{1}}=\mathbf{1}\Sigma^{-1}\boldsymbol{\alpha}$, $C={\boldsymbol{\alpha}^{\prime}\Sigma^{-1}\boldsymbol{\alpha}}$ and $D=AC-B^{2}$. All the portfolio’s are implicitly geared versions of the optimal risky portfolio $\boldsymbol{\theta}_{\alpha}$. Optimisation IV requires an explicit choice of $g_{0}$ to set the magnitude of the optimal portfolio which is typical set to one: $g_{0}=1$. The optimal portfolio controls $\boldsymbol{\theta}^{*}$ are given below each optimizations in the table. The subscript $i$ on the solution $\boldsymbol{\theta}_{i}^{*}$ is the number of the optimisation in question.
@@ 2.2 Constrained $\boldsymbol{\alpha}$-angle
$\cos^{2}(\varphi)\geq\frac{4\kappa_{\psi}}{(\kappa_{\psi}+1)^{2}}\leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ \kappa_{\psi}=\frac{\rho_{\mathrm{max}}}{\rho_{\mathrm{min}}}\frac{(1+\sin(\psi))}{(1-\sin(\psi))}.$
@@ 3 The linear gearing constraint
Table 2: Four gearing constrained portfolio optimisation problems that can be extend into the geometric framework of [19] (See Table 1). Here $A={\mathbf{1}^{\prime}\Sigma^{-1}\mathbf{1}}$,$B={\boldsymbol{\alpha}^{\prime}\Sigma^{-1}\mathbf{1}}=\mathbf{1}\Sigma^{-1}\boldsymbol{\alpha}$, $C={\boldsymbol{\alpha}^{\prime}\Sigma^{-1}\boldsymbol{\alpha}}$ and $D=AC-B^{2}$. The optimal portfolio controls $\theta^{*}$ are given below each optimisation. The GMV portfolio is $\boldsymbol{\theta}_{0}=\Sigma^{-1}\mathbf{1}/A$ and the optimal risky portfolio is $\boldsymbol{\theta}_{\alpha}=\Sigma^{-1}\boldsymbol{\alpha}/B$. See appendices E and H for the derivations for optimisations V and VIII. See sections 3.1 and 3.2 for the derivations of optimization VI and VII respectively. The subscript $i$ on the solution $\boldsymbol{\theta}_{i}^{*}$ is the number of the optimisation in question.
@@ 4.1 Constraining the $\boldsymbol{\alpha}$-angle
Table 3: Two linearly constrained portfolio optimisation problems, VI and VII, can then be extend into the robust geometric framework of [19] (See Table 1) by including a minimisation over the ball $U_{\alpha}$ defined in Equation 43. Optimization V from Table 2 is equivalent to optimization I because fixing $\sigma_{p}$ fixes the gearing $g_{0}$ (see Appendix E). Optimization IV is equivalent to VIII where the solution of IV is scaled by $g_{0}$ (see Appendix H).
@@ 5.2 Simulated $\alpha$-angle dependency
Finally, simulations were then repeated to assess whether the covariance shrinkage methodology proposed by Golts and Jones [19] for the optimal risky portfolio ($\theta_{\alpha})$ allows for consistently smaller $\alpha$-weight angles than those produced by equally-weighted portfolios, in a normally distributed environment, and in the presence of constraints. When considering constrained mean variance solutions VI and VII of Table 2, which are not simply geared versions of the optimal risky portfolio $\theta_{\alpha}$ (See Figure 5(a)) because of the linear constraints, similar results can be seen. These are evident in Figure 5(b) below. As soon as some conditioning is introduced, 100% of simulations show better $\alpha$-weight angles than those produced by the equally weighted portfolios. This confirms the theoretical implications, that only the correct formulation of conditioning creates appropriate optimal outcomes under linearly constrained quadratic optimisation. That the equally weighted portfolio, in the distribution free setting, are typically sub-optimal relative to the $\alpha$-angle shrunk covariances, and that this has been confirmed under simulation for the case of normal distributed returns.
@@
Figure 5: Both figures show the percentage number $n_{\phi_{\omega^{*}}<\phi_{1/N}}$ of performance simulations with weights better aligned with those of the returns (in circles against the left axis), and the average difference (in triangles against the right axis) between the conditioned angles and the unconditioned angles $\mu_{\Delta\phi_{k}}=\mathbb{E}_{m}[\phi_{k}-\phi_{0}]$ with confidence intervals. Fig. 5(a) is for the unconstrained optimisations, and Fig. 5(b) for the constrained cases (VI and VII). Each point represents $m=1000$ replications. This shows the unconditioned optimisations ($k$ =0) relative to those with Golts and Jones [19] covariance conditioning ($k>0$). 100% of the simulations show smaller $\alpha$-weight angles for $k>10^{-1}$ on the resulting optimal risk portfolios relative to the the equally weighted portfolio. On average, this angle is just under 30 better. Conditioned optimal portfolios outperform the equally weighted portfolios.

### v5

@@ 1 Introduction
Using a geometric construction that relates the angle between the vector of expected returns $\boldsymbol{\alpha}$ and portfolio weights $\boldsymbol{\omega}$ determined under an unconstrained mean-variance optimisation, Golts and Jones [19] were able to quantify the distortions induced by the optimisation process without making direct distributional assumptions about the underlying data generating process. They were able to link this angle directly to the condition number of the covariance matrix by showing that this angle has an upper bound dependent on the condition number. The interpretation is novel because of its geometric construction and the weak distributional assumptions this allows. When a covariance matrix is ill-conditioned, or close to degenerate, the $\boldsymbol{\alpha}$-weight angle should be very large; whereas well-conditioned matrices ensure the angle remains within tight bounds. This is much more than merely a representation change, because it provides a geometric explanation, why an equally weighted portfolio is not always preferable to a mean-variance solution even in the presence of estimation uncertainty.
@@
Reducing the angle between the weights and expected return vector is not equivalent to the implementation of a return maximisation, or return maximisation constraint. The case of pure return maximisation would result in a portfolio that weighs toward the asset with the greatest expected return, and not the direction of the expected return vector in general. In practice the $\boldsymbol{\alpha}$-weight angle should never be zero as there is always a residual correlation between the assets. In mean variance optimisations there is always a conundrum due to correlation specification error – and then determining the degree of shrinkage that should be adopted to best reflect this. The test of how much one should intervene can only be determined by out-of-period portfolio performance simulation, or with direct distributional assumptions (as demonstrated with simple simulations in Section 5).
@@ 1.1 Estimation uncertainty
Additionally, even if means and variances could accurately represent asset class distributions, the theory assumes both future expected returns, and the covariance matrix of asset returns are known, or even that they can be known. The presence of estimation error in input parameters unfortunately makes the use of Modern Portfolio Theory in its classical Markowitz formulation impractical – this is well understood and common cause [44]. The short-comings of mean-variance optimisation has been extensively discussed in the literature, with highlights including the work by Best and Grauer [6], Frankfurter et al. [16], Jobson and Korkie [29, 30], Zimmermann and Nierdermayer [50], Broadie [8], Michaud [42], and Chopra and Ziemba [12].
@@
In order to ameliorate the effect of estimation errors in the estimates of expected returns, attempts have been made to create better and more stable mean-variance optimal portfolios by utilising expected return estimators that have a better behaviour when used in the context of the mean-variance framework [44]. One of the more common techniques is the utilisation of James-Stein estimators (see Jobson and Korkie [30]) to shrink the expected returns towards the average expected return based on the volatility of the asset and the distance of its expected return from the average. Jorion [31] developed a similar technique that shrinks the expected return estimate towards the minimum variance portfolio. The area of robust statistics (see [10]) has recently been employed to create stable expected return estimates as well. Chopra and Ziemba [12] show that errors in the sample mean estimate have a larger impact on the out-of-sample performance than errors in the sample covariance estimate.
@@
There are a variety of approaches to shrinking the covariance matrix gap between the largest and smallest eigenvalues, these include optimisation based approaches [46] whose performance is measured using Sharpe ratio’s and certainty equivalents. Such approaches are novel, but do not directly address the issue of back-test over-fitting [3] nor do they provide distribution free setting with theoretical bounds for both constrained and unconstrained portfolio’s.
@@
Here we demonstrate this relationship geometrically and mathematically using an extension of the Kantorovich inequality [24, 22] provided by Bauer and Householder [4]; this elegant machinery will be shown as useful in determining why this relationship exists (outlines to both theorems are provided in A and B).
The theoretical advantage of the geometric approach is not only the distribution free setting, but the theorems directly motivate the choice of shrinkage in terms of the angle between the portfolio controls and the direction of asset returns, and thus provide a direct and unique approach. This is particularly useful if one is able to predict returns and correlations intermittently in real financial markets.
Inspiration for this idea can be seen in Axelsson [2] where much of the geometric interpretation around matrices is described.
@@ 1.2 The efficient frontier and $\alpha$-weight angle
Consider a portfolio with $n$ assets, expected return vector $\boldsymbol{\alpha}$, covariance matrix $\Sigma$, and a fixed portfolio mean $\alpha_{p}$. To find this portfolio on the efficient frontier with asset weights $\boldsymbol{\theta}$, we perform a portfolio variance minimization [40, 41]:
@@
Table 1: Some classical unconstrained portfolio optimisation problems from [19]. Here portfolios employ no explicit gearing constraint, and positions may be unconstrained long/short. Here I is a return maximisation problem with risk bound, II is a risk minimisation problem with a return bound, III is a conventional mean-variance optimisation, and IV is the Sharpe ratio optimisation. In III $\gamma$ is the inverse risk aversion parameter. In IV the risk free rate $r_{f}$ is assumed to be zero. Here $A={\mathbf{1}^{\prime}\Sigma^{-1}\mathbf{1}}$,$B={\boldsymbol{\alpha}^{\prime}\Sigma^{-1}\mathbf{1}}=\mathbf{1}\Sigma^{-1}\boldsymbol{\alpha}$, $C={\boldsymbol{\alpha}^{\prime}\Sigma^{-1}\boldsymbol{\alpha}}$ and $D=AC-B^{2}$. All the portfolio’s are implicitly geared versions of the optimal risky portfolio $\boldsymbol{\theta}_{\alpha}$. Optimisation IV requires an explicit choice of $g_{0}$ to set the magnitude of the optimal portfolio which is typical set to one: $g_{0}=1$. The optimal portfolio controls $\boldsymbol{\theta}^{*}$ are in the last column for each optimizations in the table. The subscript $i$ on the solution $\boldsymbol{\theta}_{i}^{*}$ is the number of the optimisation in question.
@@ 2.2 Constrained $\boldsymbol{\alpha}$-angle
$\cos^{2}(\varphi)\geq\frac{4\kappa_{\psi}}{(\kappa_{\psi}+1)^{2}}\leavevmode\nobreak\ \mbox{where}\leavevmode\nobreak\ \kappa_{\psi}=\frac{\rho_{\mathrm{max}}}{\rho_{\mathrm{min}}}\frac{(1+\sin(\psi))}{(1-\sin(\psi))}.$ (18)
@@ 3 The linear gearing constraint
Table 2: Four gearing constrained portfolio optimisation problems that can be extend into the geometric framework of [19] (See Table 1). Here $A={\mathbf{1}^{\prime}\Sigma^{-1}\mathbf{1}}$,$B={\boldsymbol{\alpha}^{\prime}\Sigma^{-1}\mathbf{1}}=\mathbf{1}\Sigma^{-1}\boldsymbol{\alpha}$, $C={\boldsymbol{\alpha}^{\prime}\Sigma^{-1}\boldsymbol{\alpha}}$ and $D=AC-B^{2}$. The GMV portfolio is $\boldsymbol{\theta}_{0}=\Sigma^{-1}\mathbf{1}/A$ and the optimal risky portfolio is $\boldsymbol{\theta}_{\alpha}=\Sigma^{-1}\boldsymbol{\alpha}/B$. The subscript $i$ on the solution $\boldsymbol{\theta}_{i}^{*}$ is the number of the optimisation in question.
@@ 4.1 Constraining the $\boldsymbol{\alpha}$-angle
Table 3: Two linearly constrained portfolio optimisation problems, VI and VII, can then be extend into the robust geometric framework of [19] (See Table 1) by including a minimisation over the ball $U_{\alpha}$ defined in Eqn.(44). Optimization V from Table 2 is equivalent to optimization I because fixing $\sigma_{p}$ fixes the gearing $g_{0}$. Optimization IV is equivalent to VIII where the solution of IV is scaled by $g_{0}$.
@@ 5.2 Simulated $\alpha$-angle dependency
Finally, simulations were then repeated to assess whether the covariance shrinkage methodology proposed by Golts and Jones [19] for the optimal risky portfolio ($\theta_{\alpha})$ allows for consistently smaller $\alpha$-weight angles than those produced by equally-weighted portfolios, in a normally distributed environment, and in the presence of constraints. When considering constrained mean variance solutions VI and VII of Table 2, which are not simply geared versions of the optimal risky portfolio $\theta_{\alpha}$ (Figure 5(a)) because of the linear constraints, similar results can be seen; these are evident in Figure 5(b) below. As soon as some conditioning is introduced, 100% of simulations show better $\alpha$-weight angles than those produced by the equally weighted portfolios. This implies that 100% of the simulations show improved covariance matrix conditioning, given the fact that the the matrix condition number provides an upper bound to the $\alpha$-weight angle (see equations 12 and 18).
This confirms the theoretical implications, that only the correct formulation of conditioning creates appropriate optimal outcomes under linearly constrained quadratic optimisation. That the equally weighted portfolio, in the distribution free setting, are typically sub-optimal relative to the $\alpha$-angle shrunk covariances, and that this has been confirmed under simulation for the case of normal distributed returns.
@@
Figure 5: Both figures show the percentage number $n_{\phi_{\omega^{*}}<\phi_{1/N}}$ of performance simulations with weights better aligned with those of the returns (in circles against the left axis), and the average difference (in triangles against the right axis) between the conditioned angles and the unconditioned angles $\mu_{\Delta\phi_{k}}=\mathbb{E}_{m}[\phi_{k}-\phi_{0}]$ with confidence intervals. Fig. 5(a) is for the unconstrained optimisations, and Fig. 5(b) for the constrained cases (VI and VII). Each point represents $m=1000$ replications. This shows the unconditioned optimisations ($k$ =0) relative to those with Golts and Jones [19] covariance conditioning ($k>0$). 100% of the simulations show smaller $\alpha$-weight angles for $k>10^{-1}$ on the resulting optimal risk portfolios relative to the the equally weighted portfolio. On average, this angle is just under 30 better. [19] conditioned optimal portfolios outperform the equally weighted portfolios.
5.3 Simulated Portfolio Results
Various alternatives for shrinkage estimates are available, and this section intends to briefly introduce the benefits of the Golts and Jones [19] methodology, through real data simulations. We make use of monthly total return data from all stocks included in the FTSE/JSE All Share Index from January 1996 to August 2016. This includes 332 stocks, all with varying times of listing/de-listing. A set of 1000 simulations is performed. Each simulation follows the steps outlined in Table 4.
$\alpha$-Angle Simulation Methodology
Description
1. Start date: A random start date is selected and all stocks with a 10 year history following this date are noted.
2. Investment universe; 10 stocks are randomly selected from the full set of stocks in the randomly selected 10 year window. These form the data for this simulation (120 x 10 data frame)
3. Simulation data window: The first five years of the data is used in the simulation to find the sample covariance matrix. The shrunk covariance matrix of each shrinkage methodology is then computed at each shrinkage intensity ($k$ between 0 and 1, in increments of 0.01). This five year history is also used to calculated the expected return vector, by merely using the historical mean.
4. Calculate $\theta_{\alpha}$: The optimal risk portfolio weights are found for each of the shrinkage methodologies, for varying $k$.
5. Calculate the $\alpha$-weight angle: This is found for the varying shrinkage methodologies, to see how aligned the portfolio weights are to the forward returns. The forward returns is the performance of those assets in 2 months time.
6. Out-of-sample Performance: The out-of-sample portfolio performance of the Optimal Risky portfolio, for each of the shrinkage methods, at varying shrinkage intensity ($\kappa$) is calculated. This is performed by multiplying the optimiser weights (which uses the various covariance matrices as inputs), with the performance of each asset in 2 months time. From this performance, we find the out-of-sample Sharpe ratio (SR), and a probabilistic Sharpe ratio (PSR) (see de Prado and Bailey [45]).
7. Step forward: We then roll the five year period forward one month, until 5 years (minus 1 month) of out-of-sample performance is calculated, for each of the shrinkage methodologies, at varying $k$.
Table 4: Simulation steps for the demonstration of the out-of-sample behaviour of the different shrinkage methods and their impact on the $\alpha$-weight angle. In step 5, we delay the implementation by one month and are not using the asset’s direct price 1 month later, or the return over the following months performance, because this would not be practically possible for managers to implement from a timing perspective. For practical relevance a month is skipped to allow for the real world mechanics of fund management implementation. It would be equivalent to introducing a look-a-head bias in the simulation. In addition, we note that in step 6, the probabilistic Sharpe ratio is included because it provides us with a confidence level of a particular Sharpe ratio
An important adjustment to note is that alpha-weight angles can be greater than 90 degrees (between 90 and 180 degrees) in real world data. This infrequently occurs when there are large negative expected return vectors, or large negative correlations. The Golts and Jones [19] methodology does not make provision for this and is addressed in the simulations by applying a cap on the alpha-weight angles. This means that any alpha-weight angle greater than 90 degrees, is set to the maximum value of 90 degrees.
Seven well known shrinkage methodologies are tested in this simulation. The range of shrinkage methods in the literature is significant, and this list is by no means exhaustive. For this reason we refer the interested reader to one of the many review papers given in the introduction. In terms of indicative methods we focus on those given by Ledoit [35] and Ledoit and Wolf [36, 38]. This is because we are just trying to show from this sample, the competitive nature of our proposed methodology and when we believe it should be used. The seven different methodologies we consider are given in Table 5.
Selected Shrinkage Methodologies Considered
Acr. Description
Diag. Diagonal Matrix: Shrinkage moves away from the sample covariance matrix toward the diagonal covariance matrix, $\mathrm{diag}{\Sigma}$, at intensity $k$: $\tilde{\Sigma}=(1-k)\Sigma+k\mathrm{diag}{\Sigma}$. The diagonal covariance matrix $\mathrm{diag}{\Sigma}$ is one where the diagonal of the sample covariance matrix is preserved, but the off-diagonal terms are all set to zero [35] (Appendix B2).
CCM Constant Correlation Matrix: Shrinkage moves away from the sample covariance matrix toward the constant correlation covariance matrix, $\Sigma^{CCM}$, at intensity $k$: $\tilde{\Sigma}=(1-k)\Sigma+k\Sigma^{CCM}$. Here $\Sigma^{CCM}$ the diagonal of the sample covariance matrix is preserved, but the off-diagonal terms are all set to the average of all the stock covariances [37]: $\Sigma_{ij}^{CCM}=\left\{\Sigma_{ii},i=j;\frac{1}{n(n-1)/2}\sum_{i=1}^{n}\Sigma_{ij},i\neq j\right\}.$
OPM One Parameter Matrix: Shrinkage moves away from the sample covariance matrix toward the one parameter covariance matrix, $\Sigma^{OPM}$, at intensity $k$: $\tilde{\Sigma}=(1-k)\Sigma+k\Sigma^{OPM}$. Here $\Sigma^{OPM}$ has the diagonal terms set to the the average variance, while the off-diagonal terms are all set to zero [38]: $\Sigma_{ij}^{OPM}=\left\{\frac{1}{n}\sum_{i=1}^{n}\Sigma_{ii},i=j;0,i\neq j\right\}$.
TPM Two Parameter Matrix: Shrinkage moves away from the sample covariance matrix toward the two parameter covariance matrix, $\Sigma^{TPM}$, at intensity $k$: $\tilde{\Sigma}=(1-k)\Sigma+k\Sigma^{TPM}$. Here $\Sigma^{TPM}$ is one where diagonal terms of the sample covariance matrix are all set to the the average variance, and the off-diagonal terms are all set to the average of all the stock covariances [35] (Appendix B.1): $\Sigma_{ij}^{TPM}=\left\{\frac{1}{n}\sum_{i=1}^{n}\Sigma_{ii},i=j;\frac{1}{n(n-1)/2}\sum_{i=1}^{n}\Sigma_{ij};i\neq j\right\}$.
OFMM One Factor Market Model: Shrinkage moves away from the sample covariance matrix toward the one factor market model covariance matrix, $\Sigma^{OFMM}$, at intensity $k$: $\tilde{\Sigma}=(1-k)\Sigma+k(\Sigma^{OFMM})$. Here $\Sigma^{OFMM}$ is one where the diagonal is preserved (due to the idiosyncratic volatility of the residuals), but the off-diagonal terms are shrunk to the cross-sectional average of all the random variables: $\Sigma_{ij}^{OPMM}=(s^{2}_{m}\beta\beta^{\prime}+\epsilon)_{ij}$ [36]. Here $\beta$ is the $n\times 1$ matrix of market factor loadings, $\epsilon$ a $n\times n$ diagonal matrix of residual variance estimates, and $s^{2}_{m}$ the market sample variance.
AOCS Asymptotically Optimal Convex Shrinkage: The covariance matrix is a linear shrinkage between the sample covariance matrix and the identity matrix: $\tilde{\Sigma}=(1-\alpha)\Sigma+\alpha(\tau\Sigma)$ [38].Here $\tau$ is the average sample variance and $\alpha\in[0,1]$ is the intensity parameter computed using $\alpha=\frac{1}{\mathrm{tr}([\Sigma-\tau]^{2})}\frac{1}{N}\sum_{i=1}^{N}\mathrm{tr}([z_{i}z_{i}^{T}-\Sigma]^{2})$. This is the only method investigated in this paper, where the shrinkage intensity is set (and $k$ not varied) in the simulations.
G&J Golts & Jones $\alpha$-angle minimization: The covariance matrix is a linear shrinkage between the sample covariance matrix and the identity matrix, with the $\alpha$-weight angle informing the degree of shrinkage as follows (see Equation (46)), repeated here: $\tilde{\Sigma}=\frac{k}{\cos\varphi}I+\left({1-\frac{k}{\cos\varphi}}\right)\Sigma$. Here $k$ is some quantity $\cos(\varphi_{k})$ for a targeted optimal angle $\varphi_{k}\in[\varphi,\pi/2]$ between the risky portfolio and the expected returns.
Table 5: Shrinkage method descriptions for the methods compared in the comparative portfolio simulations (Figures 6, 7 and 8).
Figure 6: The historic simulations results for the unconstrained solutions (Table 1). Fig. 6a shows the average out-of-sample Sharpe ratios, Fig. 6b shows the average out-of-sample probabilistic Sharpe ratio’s, Fig. 6c shows the average $\alpha$-weight angle, and Fig. 6d shows the average percentage improvement of the condition number of the covariance matrix; all across 1000 simulations, for varying $k$, for each shrinkage methodology given in Table 5. The G&J appears to deliver returns most aligned to future asset returns, generates higher Sharpe ratios, at high confidence levels, and improves the condition number of the covariance matrices significantly.
Figure 7: The historic simulation results for the constrained solution VI (Table 2) are shown. Fig.7a shows the average out-of-sample Sharpe ratios, Fig. 7b shows the average out-of-sample probabilistic Sharpe ratios, Fig. 7c shows the average $\alpha$-weight angle, and Fig. 7d shows the average percentage improvement of the condition number of the covariance matrix; again across 1000 simulations, for varying $k$, for each shrinkage methodology given in Table 5. While G&J methodology improves the condition number of the covariance matrices, and see’s strong $\alpha$-weight angles at small $k$, the methodology does not appear superior under this solution.
Figure 8: The historic simulation results for the constrained solution VII (Table 2) are shown. Fig . 8a shows the average out-of-sample Sharpe ratios, Fig. 8b the average out-of-sample probabilistic Sharpe ratios across 1000 simulations, for varying $k$, for each shrinkage methodology, Fig. 8c the average $\alpha$-weight angle, and Fig. 8d shows the average percentage improvement of the condition number of the covariance matrix across 1000 simulations, for varying $k$, for each shrinkage methodology given in Table 5. G&J appears to deliver returns most aligned to future asset returns,generates higher Sharpe ratios, at high confidence levels, and improves the condition number of the covariance matrices significantly.
Figure 6 provides the simulation results for the unconstrained solutions (Table 1). This includes the constrained solutions (V and VIII from Table 2) which are simply geared versions of each other. The results are shown across 1000 simulations, for varying $k$, for each shrinkage methodology. This figure shows the average out-of-sample Sharpe ratios (Figure 6a), the average out-of-sample probabilistic Sharpe ratio’s (Figure 6b), the average $\alpha$-weight angle (Figure 6c) and the average percentage improvement of the condition number of the covariance matrix (Figure 6d). Here the $\alpha$-weight angle is the angle between the ”shrunk” weights and the forward looking returns, showing which methodology will deliver performance most in line with future returns. G&J appears to deliver returns most aligned to future asset returns. We see that the G&J methodology generates higher Sharpe ratios, at high confidence levels, especially around smaller $k$ estimates, and the G&J methodology improves the condition number of the covariance matrices significantly.
Figure 7 shows the simulations results for the constrained solution VI from Table 2. The figure shows the average out-of-sample Sharpe ratios (Figure 7a), the average out-of-sample probabilistic Sharpe ratios (Figure 7b), the average $\alpha$-weight angle (Figure7c) and the average percentage improvement of the condition number of the covariance matrix (Figure 7d) across 1000 simulations, for varying $k$, for each shrinkage methodology. We see that the G&J methodology does not tend to generate higher sharpe ratios using solution VI. Again, we note that the $\alpha$-weight angle is the angle between the“shrunk” weights and the forward looking returns, showing which methodology will deliver performance most in line with future returns. While the $\alpha$-weight angle does decrease very slightly, the G&J methodology does not appear to be most favourable in this regard.
Figure 8 demonstrates the simulation results for the constrained solution VII from Table 2. Here the average out-of-sample sharpe ratios are in Figure 8a, the average out-of-sample probabilistic Sharpe ratios in Figure 8b, the average $\alpha$-weight angle in Figure 8c, and the average percentage improvement of the condition number of the covariance matrix is given in Figure 8d. The historic simulations are across 1000 simulations following the methodology outlined in Table 4, for varying $k$, and for each shrinkage methodology from Table 5. We see that the G&J methodology generates higher Sharpe ratios, at high confidence levels, especially around smaller $k$ estimates. Again, we can see which methodology will deliver performance most in line with future returns because the $\alpha$-weight angle is the angle between the “shrunk” weights and the forward looking returns. The G&J method appears to deliver returns most aligned to future asset returns, and improves the condition number of the covariance matrices significantly.
In summary, what one can conclude from the simulations shown in figures 6, 7 and 8, is that under unconstrained solutions and constrained solutions V, VII and VIII the Golts and Jones [19] method appears superior in delivering portfolios that are better aligned to the returns of the assets, and generating out-of-sample performance with superior Sharpe ratios.
However, under constrained Solution VI, this does not appear to be the case. Solution VI implements a minimum return threshold, and solutions naturally tend to solve for weights that generate that minimum return exactly. This means that all solutions are solving for the same return, and improved Sharpe ratio’s are only being generated from the standard deviation, which the G&J methodology does not necessarily solve for. Moving the portfolio weights closer to the expected return vector (shrinking the $\alpha$-weight angle) is not beneficial when the return benefit is not there.
Importantly, we tested this methodology naively using historical return means as the expected return vectors, and even with this large degree of naivety around the future asset returns, this methodology returns weights more closely resembling future asset returns. So while using historical returns is a reasonable enough forward estimate for future returns to make this method beneficial in practice, when managers have more insight, even intermittently, into future expected returns, it is clear that this methodology will be even more useful.